I have two algebra I word problems i need help setting up. I would like if you can explain the steps of setting up.
Here it is:
Problem #1
The length of a rectangle is 2 1/2 times its width. Its area is 90 square units. What are its dimensions? (Hint: length times width = area)
Let w= width in units.
L=(2 1/2)w= 5w/2 = length in units

Answer by richwmiller(17219)

now multiply length by width
(5w/2)*w=90
5w^2=45
w^2=9
w=3
5w/2=15/2=L